OpenStudy (anonymous):

solve by the elimination method 5x-7y=-53 7x+5y=59

OpenStudy (amistre64):

elimination is sooo cruel

OpenStudy (amistre64):

5x-7y=-53 [*7] 7x+5y=59 [*-5] could be a start

OpenStudy (anonymous):

amistre, which country are you from ?

OpenStudy (anonymous):

5x-7y=-53 (x5) 25x-35y=-265 7x+5y=59 (x7) 49x+35y=413 74x=148 x=2 substitute -7y=-63 y=9

OpenStudy (amistre64):

me? im from florida her ein the usa